OpenStudy (dancing_stars23):

In how many ways can a teacher arrange 6 students in the front row of a classroom with a total of 20 students?

OpenStudy (dancing_stars23):

38,760 77,520 1,860,480 27,907,200 it is either a or d, depending if it is combination or permutations

OpenStudy (dancing_stars23):

but I don't know which one

Mehek (mehek14):

\(\Large{_nC_r =} \dfrac{n!}{r!(n-r)!}\)

Mehek (mehek14):

substitute r = 6 and n = 20

Mehek (mehek14):

\(\Large{_nC_r =}\)\(\dfrac{20!}{6!(20 - 6)!}\)\(\Large{ = 38760}\)

OpenStudy (dancing_stars23):

@mehek14 thank you!